The radius of the Earth is 3.7 times the radius of the Moon, and the acceleration due to gravity on the surface of the Moon is 1/6 g, where g is the acceleration on the surface of the Earth. What is the mass of the Moon?

1
Expert's answer
2021-01-22T17:00:40-0500

The gravitational acceleration on the Moon can be written as follows:


"g_M=\dfrac{GM_M}{(R_M)^2}=\dfrac{GM_M}{(\dfrac{R_E}{3.7})^2}=13.69\dfrac{GM_M}{(R_E)^2}."

From the other hand, "g_M=\dfrac{1}{6}g_E", we can write:


"g_M=\dfrac{GM_E}{6(R_E)^2}."

We can equate both equations:


"13.69\dfrac{GM_M}{(R_E)^2}=\dfrac{GM_E}{6(R_E)^2},""13.69M_M=\dfrac{M_E}{6},""M_M=\dfrac{M_E}{13.69\cdot6}=\dfrac{M_E}{82},""M_M=\dfrac{5.97\cdot10^{24}\ kg}{82}=7.28\cdot10^{22}\ kg."